# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Abstract
Variational counterdiabatic (CD) driving is a disciplined and widely used method to robustly control quantum many-body systems by mimicking adiabatic processes with high fidelity and reduced duration. Central to this technique is a universal structure of the adiabatic gauge potential (AGP) over a parameterized Hamiltonian. Here, we reveal that introducing a new degree of freedom into the theory of the AGP can significantly improve variational CD driving. Specifically, we find that the algebraic characterization of the AGP is not unique, and we exploit this non-uniqueness to develop the weighted variational method for deriving a refined driving protocol. This approach extends the conventional method in two aspects: it effectively incorporates nonlocal information, and it assigns customized weights to matrix elements relevant to specific problems. We also develop an efficient numerical algorithm to compute the refined driving protocol using computer algebra. Our framework is broadly applicable, as it can replace almost all previous uses of variational CD driving. We demonstrate its practicality by applying it to adiabatic evolution along the ground state of a parameterized Hamiltonian. This proposal outperforms the conventional method in terms of fidelity, as confirmed by extensive numerical simulations on quantum Ising models.
@@ Introduction
Fast, accurate, and stable control of quantum systems is of paramount importance to realize quantum state preparation, quantum state processing, and quantum information processing for advancing quantum technologies [1]. High-quality quantum control is realized by combining a theoretical scheme to compute a driving protocol on classical computers with experimental techniques to faithfully implement the protocol, both forming broad research areas [2]. While the early development of quantum control focused on few-body systems, recent advances in quantum technology highlight the need for methods tailored to many-body systems, such as systems with tens to hundreds of qubits [3, 4, 5].
@@
Despite active research, controlling many-body systems remains a huge theoretical challenge. One of the fundamental difficulties is the exponential computational cost of simulating quantum dynamics on classical computers. Control methods such as quantum optimal control [2] and reinforcement learning [6] rely on time-evolution simulations to maximize the final fidelity, but direct simulation requires exponentially large memory and time for many-body systems. Compression methods such as tensor networks have been incorporated to reduce this cost [7], but they involve tradeoffs between accuracy and computational cost. In particular, this reliance on classical simulation is inappropriate for quantum computation, which should exploit quantum dynamics beyond classical simulations.
@@
An alternative approach that avoids time-evolution simulations is to determine driving protocols from fundamental and universal structures of quantum systems. One such structure is the adiabatic gauge potential (AGP), which endows a geometric structure with a parameterized Hamiltonian [8]. By implementing the AGP as a driving Hamiltonian, we can perform counterdiabatic (CD) driving [9, 10, 11, 12]. CD driving mimics adiabatic processes with—in principle—an arbitrarily short duration, with the increased energy cost manageable by the quantum speed limit [13, 14, 15, 16]. CD driving inherits the robustness of adiabatic processes [17, 18] and is widely used to stably guide a quantum state along an eigenstate of a time-dependent Hamiltonian. Explicit expressions of the AGP have been derived for a variety of simple systems, including two- and three-level systems [19], harmonic oscillators [20], many-body quantum spin systems associated with free fermions [21], classical particles in potentials [22], many-body and/or nonlinear systems with scale-invariant dynamics [23, 24], quantum systems associated with classical nonlinear integrable systems [25], and classical spin systems [26].
@@
In recent years, the variational AGP, a systematic approximation of the AGP [27, 8], has attracted attention for its ability to overcome the difficulty of computing and implementing the exact AGP in many-body systems. The variational AGP is defined by minimizing an action (a functional of operators) over experimentally feasible operators. The action is derived from an algebraic characterization of the exact AGP [22], and it could yield the exact AGP if minimized over arbitrary Hermitian operators. The minimization can be easily computed even for large systems, making the variational AGP practical for many-body systems. Using the variational AGP as a driving Hamiltonian, we can realize approximated CD driving, which is called variational CD driving or local CD driving.
@@
Variational CD driving has been tested and applied to a wide range of systems, leading to significantly improved quantum control. Numerical tests have been performed on spin systems with various lattice structures and interactions [28, 29, 30, 31, 32, 33] and on fermion lattice systems [34, 27]. It has been realized experimentally with nuclear magnetic resonance [35], a tight-binding lattice of ultracold atoms [36], and superconducting quantum computers [37]. Practical applications have been considered and tested, such as heat engines [38, 39], quantum state transfer [40, 41], quantum annealing and its varieties [28, 29, 30, 31, 32, 33, 42], and optimization problems including integer factorization, portfolio optimization, protein folding, and logistics [43, 44, 45, 46, 47, 48]. In addition, variational CD driving has been combined with other techniques and ideas to derive a number of practical methods, including Floquet engineering [49, 50], tensor networks [51, 52], digital quantum simulation [43, 44], Lie algebra [53], the Lanczos algorithm [54, 55], and various optimizations of the intermediate path [31, 56, 57]. From a broader perspective, variational AGPs have been extended to variational quantum algorithms [58, 59, 60, 61], approximate diagonalization [62, 63], open quantum systems [64, 65], and classical Hamiltonian systems [66], and they have been used as a sensitive probe for quantum phase transitions [53, 55], quantum chaos [54], and singular macroscopic degeneracies [67]. In most applications, variational CD driving significantly improves the driving quality, such as the fidelity to the target state, compared to the bare time evolution without adding the variational AGP. Nevertheless, the fidelity often remains far from unity for complex systems, suggesting that the existing variational CD driving is not in its optimal form.
@@
In fact, the existing variational AGP leaves one fundamental aspect unaddressed. The variational method has always used a unique form of action. However, any functionals that attain their minimum at the exact AGP could, in principle, play the role of action for determining a variational AGP. There should be an infinite number of such functionals. Different actions should yield different approximated AGPs and may provide improved quantum control. Although some literature mentions the possibility of constructing an alternative action using a Gibbs state [27, 55], this action is difficult to compute in many-body systems without compromising practicality.
@@
In this paper, we address this gap by both advancing the fundamental theory of the AGP and developing a computational technique, revealing that the conventional approach did not fully exploit the potential of variational CD driving. We find that the algebraic characterization of the AGP is not unique, and we introduce an infinite variety of characterizations. By exploiting this freedom, we develop a systematic framework for constructing actions tailored to specific purposes, termed weighted actions. The weighted action extends the conventional approach in two aspects. First, it allows us to assign weights to relevant energy eigenstates to reduce the approximation error of important matrix elements. Second, the weighted action contains nonlocal terms, allowing the resulting variational AGPs to reflect nonlocal information of the system effectively, in contrast to the conventional action consisting of local terms.
@@
We show that this theoretical advance can be practically implemented by using computer algebra in general spin and fermionic systems. Computer algebra is the computation performed with algebraic relations between operators rather than matrix representations [68], and it can compute variational AGPs from weighted actions in polynomial time with respect to the system size. In particular, we construct a concrete algorithm for general spin-1/2 systems and provide an explicit computational time analysis, extending similar algebraic methods implicitly used in the literature to compute AGPs [34, 33]. An efficient C++ implementation is available on GitHub [69].
@@
This proposal, which we term the weighted variational method, is practical and widely applicable, as it can—in principle—enhance almost all previous applications of variational AGPs by simply replacing the conventional action with weighted actions. This enhancement is experimentally implementable without substantial new difficulties such as new driving fields/interactions or faster control of the driving fields. To demonstrate the power of our proposal, we apply it to quantum control along the ground state of an arbitrary time-dependent Hamiltonian by designing a weighted action that emphasizes lower-energy eigenstates. We numerically test this application in quantum Ising models with random couplings. The final ground-state fidelity increases in almost all samples, and the median increase reaches, for example, 19 times for spin-glass systems. This enhancement is achieved even in the short-duration limit and remains valid for larger system sizes.
@@
The remainder of this paper is organized as follows. In Secs. IIIV, we present our proposal at three different levels of generality. In Sec. II, we develop the most general framework, including the extended algebraic characterizations of the AGP, weighted actions, and the numerical algorithm with computer algebra. Section III is at the second level of generality, where we specialize the framework to ground-state evolution. This specialization is still valid for arbitrary quantum systems. Section IV presents the most concrete results obtained from numerical simulations of quantum Ising models. Section V discusses the possible mechanisms behind this improvement using numerical analysis and the quantum speed limit formula. Section VI concludes the paper.
@@
Setup and preliminaries
We start with a general setup: consider a general quantum system—few-body or many-body—with a parameterized Hamiltonian $H(\lambda)$. Without loss of generality, we assume that $\lambda$ is scalar and takes $0\leq\lambda\leq 1$. We aim to realize an adiabatic process from $\lambda=0$ to $\lambda=1$, which keeps the population of each eigenstate intact except at degeneracies. If we are allowed to take a very long time, this goal is achieved simply by driving the system with the Hamiltonian $H(\lambda)$ with slowly modulating $\lambda$ from 0 to 1, as ensured by the adiabatic theorem [70]. However, practical driving protocols must achieve the process in a short time to combat decoherence and dissipation.
A quick adiabatic process can be achieved by CD driving. We fix an arbitrary schedule $\lambda_{t}$ and consider adding a driving Hamiltonian of the form $\dot{\lambda}_{t}V(\lambda_{t})$, where the dot denotes the time derivative, and $V(\lambda)$ is a $\lambda$-dependent driving Hamiltonian. We evolve the system according to the Schrödinger equation,
@@
$H(\lambda)=\sum_{n=1}^{D}\epsilon_{n}(\lambda)|\phi_{n}(\lambda)\rangle\langle\phi_{n}(\lambda)|,$ (2)
@@
Here, $\kappa_{n}(t)=\epsilon_{n}(\lambda_{t})+\dot{\lambda}_{t}[V(\lambda_{t})]_{nn}-i\hbar\dot{\lambda}_{t}\langle\phi_{n}(\lambda_{t})|\partial_{\lambda}\phi_{n}(\lambda_{t})\rangle$ is a real-valued function, $[\cdots]_{nm}$ is the shorthand for $\langle\phi_{n}(\lambda_{t})|\cdots|\phi_{m}(\lambda_{t})\rangle$, and $\Phi(\lambda)$ is a Hermitian operator called the AGP,
$\Phi(\lambda)\coloneqq\sum_{\begin{subarray}{c}nm\\ (n\neq m)\end{subarray}}i\hbar|\phi_{m}(\lambda)\rangle\langle\phi_{m}(\lambda)|\partial_{\lambda}\phi_{n}(\lambda)\rangle\langle\phi_{n}(\lambda)|.$ (4)
In the right-hand side of Eq. (3), the first term $\kappa_{n}(t)c_{n}(t)$ only changes the complex phase of $c_{n}(t)$, keeping the population $|c_{n}(t)|^{2}$ intact, but the second term causes undesired nonadiabatic transitions, i.e., transitions from one eigenstate to another. The nonadiabatic transitions can be perfectly suppressed by choosing $V(\lambda)=\Phi(\lambda)$, which is called CD driving [9, 10, 11, 12].
The AGP is written in terms of the eigenstates of the Hamiltonian, which poses two difficulties for CD driving of many-body systems. First, the calculation of the AGP involves the diagonalization of the Hamiltonian $H(\lambda)$, which is usually impossible for many-body systems. Second, the AGP contains nonlocal many-body interactions, which hinders experimental implementation.
To circumvent these two difficulties, Ref. [27] proposed a variational method to determine approximated AGPs. The variational method starts with an algebraic characterization of the exact AGP [22],
$[H(\lambda),\partial_{\lambda}H(\lambda)-i\hbar^{-1}[H(\lambda),\Phi(\lambda)]]=0.$ (5)
@@
where $\|A\|\coloneqq\sqrt{\mathop{\mathrm{tr}}(A^{\dagger}A)}$ denotes the Hilbert–Schmidt norm. The minimization is performed over any Hermitian operator $V$, and $\in$” indicates that the minimizer of $\mathcal{S}_{\lambda}^{(1)}[V]$ is not unique, and $\Phi(\lambda)$ is one of them. The meaning of the superscript “$(1)$” will be clarified later.
The variational method determines the variational AGP by optimizing the action $\mathcal{S}^{(1)}_{\lambda}[V]$ over a set of experimentally implementable Hermitian operators. For concreteness, we consider an ansatz of the form
@@
where $\bm{\alpha}(\lambda)\equiv(\alpha_{1}(\lambda),\ldots,\alpha_{M}(\lambda))$ are $\lambda$-dependent scalar coefficients to be optimized, and $A_{1},\dots,A_{M}$ are $\lambda$-independent Hermitian operators. For practical applications, $A_{\mu}$’s are taken to be experimentally feasible operators. For each $\lambda$, the coefficient $\bm{\alpha}(\lambda)$ is determined by the following minimization:
@@
where “const.” refers to a term independent of $V$. This expression shows that the minimization of $\mathcal{S}_{\lambda}^{(1)}[V]$ makes $V_{mn}$ as close to $[\Phi(\lambda)]_{mn}$ as possible under a given constraint.
@@
While the variational method has proven useful for a wide range of systems, it may not provide the best possible driving protocol, as suggested by the following two arguments. First, Eq. (9) shows that the action $\mathcal{S}_{\lambda}^{(1)}[V]$ is the sum over $|[V-\Phi(\lambda)]_{mn}|^{2}$ for all $(m,n)$, and the minimization of $\mathcal{S}_{\lambda}^{(1)}[V]$ tries to suppress all of these $O(D^{2})$ terms. On the other hand, the minimization is conducted with only $M$ degrees of freedom, which is usually much smaller than $O(D^{2})$. Thus, the minimization problem is severely underparametrized, and suppressing one element may conflict with suppressing another. This fact suggests that we can improve variational CD driving by introducing an alternative action that assigns weights to relevant elements, prioritizing the suppression of errors in important elements over unimportant ones.
Second, the action $\mathcal{S}_{\lambda}^{(1)}[V]$ often fails to incorporate nonlocal information into determining variational AGPs. In simple systems, the coefficient $\alpha_{\mu}^{(1)}(\lambda)$ from Eq. (8) often depends only on local parameters. For example, in the Ising model with $A_{\mu}$ acting on one spin, the coefficient $\alpha_{\mu}(\lambda)$ is determined solely by the local magnetic field and the nearest-neighbor coupling constants around the spin [27]. More generally, the action $\mathcal{S}_{\lambda}^{(1)}[V]$ is always a sum of local contributions whenever $H(\lambda)$ and $A_{\mu}$’s are sums of local operators (see Appendix B.1). This locality is unfavorable for approximating the exact AGP, which is usually nonlocal and sensitive to small perturbations in complex many-body systems [71]. Therefore, variational CD driving may be improved by using an alternative action that incorporates nonlocal information.
Motivated by these arguments, we introduce a systematic framework for constructing alternative actions, termed weighted actions. Alternative actions should solve the above two issues of the conventional action. At the same time, they should be easy to compute using the algebraic relations between operators without diagonalizing the Hamiltonian.
@@
$[H(\lambda),\partial^{\prime}_{\lambda}\mathcal{P}_{\lambda}(H(\lambda))-i\hbar^{-1}[\mathcal{P}_{\lambda}(H(\lambda)),\Phi(\lambda)]]=0.$ (11)
@@
This expression extends Eq. (6). When $\mathcal{P}(x)$ is a degree one polynomial ($K=1$), the action $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$ is equivalent to the conventional action $\mathcal{S}_{\lambda}^{(1)}[V]$ except for an overall multiplicative factor, which follows from $\partial^{\prime}_{\lambda}[p_{1}(\lambda)H(\lambda)]=p_{1}(\lambda)\partial_{\lambda}H(\lambda)$ and $\partial^{\prime}_{\lambda}p_{0}(\lambda)=0$.

### v3

@@ Abstract
Variational counterdiabatic (CD) driving is a disciplined and widely used method to robustly control quantum many-body systems by mimicking adiabatic processes with high fidelity and reduced duration. Central to this technique is a universal structure of the adiabatic gauge potential (AGP) over a parameterized Hamiltonian. Here, we reveal that introducing a new degree of freedom into the theory of the AGP can significantly improve variational CD driving. Specifically, we find that the algebraic characterization of the AGP is not unique, and we exploit this non-uniqueness to develop the weighted variational method for deriving a refined driving protocol. This approach extends the conventional method in two aspects: it assigns customized weights to matrix elements relevant to specific problems, and it effectively incorporates nonlocal information. We also develop an efficient numerical algorithm to compute the refined driving protocol using computer algebra. Our framework is broadly applicable and, in principle, it can replace any previous use of variational CD driving. We demonstrate its practicality by applying it to adiabatic evolution along the ground state of a parameterized Hamiltonian. This proposal outperforms the conventional method in terms of fidelity, as confirmed by extensive numerical simulations on quantum Ising models.
@@ Introduction
Fast, accurate, and stable control of quantum systems is of paramount importance to realize quantum state preparation, state processing, and information processing for advancing quantum technologies [1]. High-quality quantum control is realized by combining a theoretical scheme to compute a driving protocol on classical computers with experimental techniques to faithfully implement the protocol, both forming broad research areas [2]. While the early development of quantum control focused on few-body systems, recent advances in quantum technology highlight the need for methods tailored to many-body systems, such as systems with tens to hundreds of qubits [3, 4, 5].
@@
Controlling many-body systems poses a huge theoretical challenge, especially due to the exponential computational cost of simulating quantum dynamics on classical computers. Control methods such as quantum optimal control [2] and reinforcement learning [6] rely on time-evolution simulations to maximize the final fidelity, but direct simulation requires exponentially large memory and time for many-body systems. Compression methods such as tensor networks have been incorporated to reduce this cost [7], but they involve tradeoffs between accuracy and computational cost. In particular, this reliance on classical simulation is inappropriate for quantum computation, which should exploit quantum dynamics beyond classical simulations.
@@
An alternative approach that avoids time-evolution simulations is to determine driving protocols from fundamental and universal structures of quantum systems. One such structure is represented by an operator called the adiabatic gauge potential (AGP), which endows a geometric structure with a parameterized Hamiltonian [8]. By implementing the AGP as a driving Hamiltonian, we can perform counterdiabatic (CD) driving [9, 10, 11, 12]. CD driving mimics adiabatic processes with—in principle—an arbitrarily short duration, with the increased energy cost manageable by the quantum speed limit [13, 14, 15, 16]. CD driving inherits the robustness of adiabatic processes [17, 18] and is widely used to stably guide a quantum state along an eigenstate of a time-dependent Hamiltonian. Explicit expressions of the AGP have been derived for a variety of simple systems, including two- and three-level systems [19], harmonic oscillators [20], many-body quantum spin systems associated with free fermions [21], classical particles in potentials [22], many-body and/or nonlinear systems with scale-invariant dynamics [23, 24], quantum systems associated with classical nonlinear integrable systems [25], and classical spin systems [26].
@@
In recent years, the variational AGP, a systematic approximation of the AGP [27, 8], has attracted attention for its ability to overcome the difficulty of computing and implementing the exact AGP in many-body systems. The variational AGP is defined by minimizing an action (a functional of operators) over experimentally feasible operators. The action is derived from an algebraic characterization of the exact AGP [22], and it could yield the exact AGP if minimized over arbitrary Hermitian operators. The minimization can be easily computed even for large systems, making the variational AGP practical for many-body systems. Using the variational AGP as a driving Hamiltonian, we can realize approximated CD driving, called variational CD driving or local CD driving.
@@
Variational CD driving has been tested and applied to a wide range of systems, leading to significantly improved quantum control. Numerical tests have been performed on spin systems with various lattice structures and interactions [28, 29, 30, 31, 32, 33, 34] and on fermion lattice systems [35, 27]. It has been experimentally realized with nuclear magnetic resonance [36], a tight-binding lattice of ultracold atoms [37], and superconducting quantum computers [38]. Practical applications have been considered and tested, such as heat engines [39, 40], quantum state transfer [41, 42], quantum annealing and its varieties [28, 29, 30, 31, 32, 33, 43], and optimization problems including integer factorization, portfolio optimization, protein folding, and logistics [44, 45, 46, 47, 48, 49]. In addition, variational CD driving has been combined with other techniques and ideas to derive a number of practical methods, including Floquet engineering [50, 51, 52, 53, 54], tensor networks [55, 56], digital quantum simulation [44, 45], Lie algebra [57], the Lanczos algorithm [58, 59], and optimizations of the intermediate path [31, 60, 61]. From a broader perspective, variational AGPs have been extended to variational quantum algorithms [62, 63, 64, 65], approximate diagonalization [66, 67], open quantum systems [68, 69], and classical Hamiltonian systems [70], and they have been used as a sensitive probe for quantum phase transitions [57, 59], quantum chaos [58], and singular macroscopic degeneracies [71].
In most applications, variational CD driving significantly improves the driving quality, such as the fidelity to the target state, compared to the bare time evolution without adding the variational AGP. Nevertheless, the fidelity often remains far from unity, particularly for many-body and random systems (e.g., Ref. [29]). This is presumably due to two key insufficiencies in the driving protocols determined by variational CD driving. First, the driving protocol is ignorant of which eigenstate is targeted, for example, whether it is the ground state or the highest-energy state. Second, the local driving coefficient of the protocol is often determined by the system’s local parameters [27, 28] and fails to globally integrate nonlocal information, which is unfavorable for controlling highly entangled eigenstates of many-body systems.
@@
To overcome these two difficulties, this paper focuses on a fundamental yet unaddressed aspect of variational AGP. The variational method has always employed a specific form of action. However, any functionals that attain their minimum at the exact AGP could, in principle, play the role of action for determining a variational AGP. There should be an infinite number of such functionals. Different actions should yield different approximated AGPs and may help overcome the key issues of the existing method. Although some literature mentions an alternative action involving a Gibbs state [27, 59], this action is difficult to compute in many-body systems without compromising practicality.
@@
In this paper, we materialize this observation by both advancing the fundamental theory of the AGP and developing a computational technique, revealing that the conventional approach did not fully exploit the potential of variational CD driving. We find that the algebraic characterization of the AGP is not unique, and we introduce an infinite variety of characterizations. By exploiting this freedom, we develop a systematic framework for constructing actions tailored to specific purposes, termed weighted actions. The weighted action overcomes the above two insufficiencies of the conventional method. First, it allows us to assign weights to relevant energy eigenstates to reduce the approximation error of important matrix elements. Second, the weighted action contains nonlocal terms, allowing the resulting variational AGPs to reflect nonlocal information of the system effectively.
@@
We show that this theoretical advance can be practically implemented by using computer algebra in general spin and fermionic systems. Computer algebra is the computation performed with algebraic relations between operators rather than matrix representations [72], and it can compute variational AGPs from weighted actions in polynomial time with respect to the system size. In particular, we construct a concrete algorithm for general spin-$1/2$ systems and provide an explicit computational time analysis, extending similar algebraic methods implicitly used in some literature to compute AGPs [35, 33]. An efficient C++ implementation is available in Ref. [73].
@@
This proposal, which we term the weighted variational method, is practical and widely applicable, as it can—in principle—generalize any previous application of variational AGPs by simply replacing the conventional action with weighted actions. This generalization is experimentally implementable without substantial new difficulties, such as new driving fields/interactions or faster control of the driving fields. To demonstrate the power of our proposal, we apply it to quantum control along the ground state of an arbitrary time-dependent Hamiltonian by designing a weighted action that emphasizes lower-energy eigenstates. We numerically test this application in quantum Ising models with random couplings. The final ground-state fidelity increases in almost all samples, and the median increase reaches, for example, 19 times for spin-glass systems. This enhancement is achieved even in the short-duration limit and remains valid for larger system sizes.
@@
The remainder of this paper is organized as follows. In Secs. IIIV, we present our proposal at three different levels of generality. In Sec. II, we develop the most general framework, including the extended algebraic characterizations of the AGP, weighted actions, and the numerical algorithm with computer algebra. Section III is at the second level of generality, where we specialize the framework to ground-state evolution. This specialization is still valid for arbitrary quantum systems. Section IV presents the most concrete results obtained from numerical simulations of quantum Ising models. Section V analyzes the possible mechanisms behind the improvement in fidelity using numerical analysis and the quantum speed limit formula. Section VI concludes the paper.
@@
Setup
We start with a general setup: consider a general quantum system—few-body or many-body—with a parameterized Hamiltonian $H(\lambda)$. Without loss of generality, we assume that $\lambda$ is scalar and takes values $0\leq\lambda\leq 1$. We aim to realize an adiabatic process from $\lambda=0$ to $\lambda=1$, keeping the population of each eigenstate intact except at degeneracies. If we are allowed to take a very long time, this goal is achieved simply by driving the system with the Hamiltonian $H(\lambda)$ with slowly modulating $\lambda$ from 0 to 1, as ensured by the adiabatic theorem [74]. However, practical driving protocols must achieve the process in a short time to combat decoherence and dissipation.
A quick adiabatic process with an arbitrary schedule $\lambda_{t}$ can be achieved by CD driving. Consider adding a driving Hamiltonian of the form $\dot{\lambda}_{t}V(\lambda_{t})$, where the dot denotes the time derivative, and $V(\lambda)$ is an operator depending on $\lambda$. We evolve the system according to the Schrödinger equation,
@@
$H(\lambda)=\sum_{n=1}^{D}\epsilon_{n}(\lambda)|\phi_{n}(\lambda)\rangle\langle\phi_{n}(\lambda)|$ (2)
@@
Here, $\kappa_{n}(t)\coloneqq\epsilon_{n}(\lambda_{t})+\dot{\lambda}_{t}[V(\lambda_{t})]_{nn}-i\hbar\dot{\lambda}_{t}\langle\phi_{n}(\lambda_{t})|\partial_{\lambda}\phi_{n}(\lambda_{t})\rangle$ is a real-valued function, $\cdots]_{nm}$ is the shorthand for $\langle\phi_{n}(\lambda_{t})|\cdots|\phi_{m}(\lambda_{t})\rangle$, and $\Phi(\lambda)$ is a Hermitian operator called the AGP,
$\Phi(\lambda)\coloneqq i\hbar\sum_{\begin{subarray}{c}nm\\ (n\neq m)\end{subarray}}|\phi_{n}(\lambda)\rangle\langle\phi_{n}(\lambda)|\partial_{\lambda}\phi_{m}(\lambda)\rangle\langle\phi_{m}(\lambda)|.$ (4)
In the right-hand side of Eq. (3), the first term $\kappa_{n}(t)c_{n}(t)$ only changes the complex phase of $c_{n}(t)$, keeping the population $|c_{n}(t)|^{2}$ intact, but the second term causes undesired nonadiabatic transitions, i.e., transitions from one eigenstate to another. The nonadiabatic transitions can be perfectly suppressed by choosing $V(\lambda)=\Phi(\lambda)$, which is called CD driving [9, 10, 11, 12]. In the energy eigenbasis of $H(\lambda)$, the CD driving term $V(\lambda)$ is intuitively depicted as a “hammer” that exerts an external force on the system to make it follow an eigenstate faithfully [Fig. 1(a)].
The exact CD driving, however, has two crucial drawbacks when applied to many-body systems. First, the calculation of the AGP involves the diagonalization of the Hamiltonian $H(\lambda)$, which is usually impossible for many-body systems. Second, the nonlocal many-body interactions in the AGP hinder experimental implementation. Implementable control fields are usually local, such as one- and two-body. It has been proposed to use Floquet engineering to effectively expand the implementable set of control fields, including nonlocal interactions [50, 51, 52, 53], but the number of independently controllable fields is still practically limited due to the constraint on the highest achievable control frequency. It thus remains difficult to implement the full AGP in many-body systems.
Figure 1: Schematics of the methods of CD driving. (a) Concept of exact CD driving, depicted in the energy eigenbasis of the original Hamiltonian $H(\lambda)$. Gray curves represent the energy of the eigenstates of $H(\lambda)$. Without CD driving, the population escapes from the target eigenstates (e.g., the ground state in this figure) during the time evolution, as depicted by the red curve (left). CD driving uses a driving force $V(\lambda)=\Phi(\lambda)$, depicted as a “hammer” hitting the system, to ensure that the system correctly tracks the target eigenstate during time evolution (right). Exact CD driving keeps the system perfectly on the target, but it is impractical for many-body systems. (b) Comparison between the conventional and weighted variational methods of CD driving. The conventional variational method tries to design a hammer $V(\bm{\alpha}(\lambda))$ that hits all eigenstates equally well (left). However, this goal may not be fully achieved with a limited number of driving fields. In contrast, our weighted variational method designs a specialized hammer for important eigenstates (e.g., low-energy states in this figure), providing a tailored approach to a specific problem (right). As a result, our method can more effectively suppress the escape of the population from the target eigenstates, as shown by the red curves.
Conventional variational method
Given these practical difficulties in implementing the full AGP, it is desirable to establish an optimal approximation of CD driving within the limited control fields at hand. Reference [27] addressed this problem by proposing a variational method to determine approximated AGPs.
The variational method starts with an algebraic characterization of the exact AGP [22],
$H(\lambda),\partial_{\lambda}H(\lambda)-i\hbar^{-1}[H(\lambda),\Phi(\lambda)]]=0.$ (5)
@@
where $\|A\|\coloneqq\sqrt{\mathop{\mathrm{tr}}(A^{\dagger}A)}$ denotes the Hilbert–Schmidt norm. The minimization is performed over any Hermitian operator $V$. Note that this variational formula only fixes the part of $V$ non-commuting with $H(\lambda)$, as the action $\mathcal{S}_{\lambda}^{(1)}[V]$ is independent of the part of $V$ that commutes with $H(\lambda)$. Thus, the minimizer of $\mathcal{S}_{\lambda}^{(1)}[V]$ is not unique, as indicated by the symbol $\in$”, and $\Phi(\lambda)$ is one of them. The meaning of the superscript “(1)(1)” will be clarified later. Since the action $\mathcal{S}_{\lambda}^{(1)}[V]$ is minimized by the true AGP, the difference $\mathcal{S}_{\lambda}^{(1)}[V]-\mathcal{S}_{\lambda}^{(1)}[\Phi(\lambda)]$ can be interpreted as quantifying how far $V$ is apart from $\Phi(\lambda)$ without explicitly using the matrix elements of $V$ and $\Phi(\lambda)$.
Motivated by this interpretation, the variational method determines the variational AGP by minimizing the action $\mathcal{S}^{(1)}_{\lambda}[V]$ over an ansatz consisting of experimentally implementable Hermitian operators. For concreteness, we consider an ansatz of the form
@@
where $\bm{\alpha}(\lambda)\equiv(\alpha_{1}(\lambda),\ldots,\alpha_{M}(\lambda))$ are $\lambda$-dependent scalar coefficients to be optimized, and $A_{1},\dots,A_{M}$ are $\lambda$-independent Hermitian operators. For practical applications, $A_{\mu}$’s are taken to be experimentally feasible operators. For each $\lambda$, the coefficient $\bm{\alpha}(\lambda)$ is determined by the minimization,
@@
where “const.” refers to a term independent of $V$. This expression shows that the minimization of $\mathcal{S}_{\lambda}^{(1)}[V]$ makes $V_{mn}$ as close to $\Phi(\lambda)]_{mn}$ as possible within a given ansatz.
@@
Although the driving coefficient $\bm{\alpha}^{(1)}(\lambda)$ is determined through an optimization, it is not guaranteed to provide the best possible driving protocol because the optimization depends on the choice of variational functional, i.e., how we measure the “distance” from the exact AGP. In fact, the conventional variational functional $\mathcal{S}^{(1)}_{\lambda}$ has two significant limitations, despite its widespread applications.
First, the conventional action does not reflect which eigenstate is the target of the driving. Equation (9) shows that the action $\mathcal{S}_{\lambda}^{(1)}[V]$ is the sum over the matrix elements $|[V-\Phi(\lambda)]_{mn}|^{2}$ for all $(m,n)$, and the minimization of $\mathcal{S}_{\lambda}^{(1)}[V]$ tries to suppress all of these $O(D^{2})$ terms. On the other hand, minimization is carried out with only $M$ degrees of freedom, which is usually much smaller than $O(D^{2})$. Thus, the minimization problem is severely underparameterized, and suppressing one matrix element may conflict with suppressing another. This fact suggests that we can improve variational CD driving by introducing an alternative action that assigns weights to relevant energy eigenstates, prioritizing the suppression of errors in important elements over unimportant ones.
Second, the action $\mathcal{S}_{\lambda}^{(1)}[V]$ often fails to incorporate nonlocal information into variational AGPs. In simple systems, the coefficient $\alpha_{\mu}^{(1)}(\lambda)$ from Eq. (8) often depends only on local parameters. For example, in the Ising model with $A_{\mu}$ acting on one spin, the coefficient $\alpha_{\mu}(\lambda)$ is determined solely by the local magnetic field and the nearest-neighbor coupling constants around the spin [27]. More generally, the action $\mathcal{S}_{\lambda}^{(1)}[V(\bm{\alpha})]$ is always a sum of local contributions whenever $H(\lambda)$ and $A_{\mu}$’s are sums of local operators (see Appendix A.1). This locality is unfavorable for approximating the exact AGP, which is usually nonlocal and sensitive to small perturbations in complex many-body systems [75]. Therefore, variational CD driving may be improved by using an alternative action that incorporates nonlocal information.
Motivated by these arguments, we introduce a systematic framework for constructing alternative actions, termed weighted actions. Alternative actions should solve the above two issues of the conventional action. At the same time, they should be easy to compute without diagonalizing the Hamiltonian.
@@
$H(\lambda),\partial^{\prime}_{\lambda}\mathcal{P}_{\lambda}(H(\lambda))-i\hbar^{-1}[\mathcal{P}_{\lambda}(H(\lambda)),\Phi(\lambda)]]=0.$ (11)
@@
This expression extends Eq. (6), and the minimizer is unique up to the part of $V$ that commutes with $\mathcal{P}_{\lambda}(H(\lambda))$. When $\mathcal{P}_{\lambda}(x)$ is a degree one polynomial ($K=1$), the action $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$ is equivalent to the conventional action $\mathcal{S}_{\lambda}^{(1)}[V]$ except for an overall multiplicative factor, which follows from $\partial^{\prime}_{\lambda}[p_{1}(\lambda)H(\lambda)]=p_{1}(\lambda)\partial_{\lambda}H(\lambda)$ and $\partial^{\prime}_{\lambda}p_{0}(\lambda)=0$. The difference $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]-\mathcal{S}_{\lambda}^{(\mathcal{P})}[\Phi(\lambda)]$ of this action can be interpreted as measuring the “distance” between $V$ and $\Phi(\lambda)$ in a way different from the original action and dependent on the choice of the polynomial. This distance measure is still written without using the matrix elements of $V$ and $\Phi(\lambda)$.